OpenStudy (anonymous):

If I have the integral r^4lnrdr, should it be solved using parts?

OpenStudy (anonymous):

\[ \int\limits_{1}^{3} r^4lnr dr\]

OpenStudy (anonymous):

Ohhhh

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

Yes....u = r^4 v' = ln r

OpenStudy (anonymous):

Yes, U=r^4 dU=4r^3 v=x*ln(x)-x dv=ln(x)

OpenStudy (anonymous):

Sorry, is with r xD... U=r^4 dU=4r^3 v=r*ln(r)-r dv=ln(r)

OpenStudy (anonymous):

\[U=r^4 dU=4r^3 v=r*\ln(r)-r dv=\ln(r)\]

OpenStudy (anonymous):

\[\int\limits r^4 \log (r) dr = 1/25*R^5 (5\log (r) -1) + c\]

OpenStudy (anonymous):

For the above solution....you'd have to know the integral for ln(r), which is not a common integral. I prefer u=ln(r), dv=r^4dr....work from there.

OpenStudy (anonymous):

Yeah, I goofed up, u is supposed to be ln(r)